# Adversarial Reasoning for Repair Based on Inferred Program Intent

### v1

@@ 3.2. Agenttest
First, Agenttest generates tests to measure the adversarial degree of inferred intents. If the adversarial degree of an intent is low, that intent will be removed, and a new adversarial intent will be generated. Second, Agenttest generates additional tests to address the overfitting problem by validating the patches generated by Agentrepair. Specifically, we validate whether the generated patches are overfitting to a specific intent. When one of the intents is close to the developer’s intent, the generated patches are considered correct rather than overfitting. Note that this use of tests generated as an integral part of repair is, to our knowledge, novel. Overfitting patch assessment tends to be considered a post-processing step in patch generation, or left to manual effort (as in previous agent-based approaches (repairagent, ; fix-agent, ; autocoderover, ), which conclude repair attempts upon finding plausible test-passing patches).
@@
Recall that Agentreason produces multiple adversarial program intents. Similarly, Agenttest aims to generate adversarial tests for these inferred intents. In this context, adversarial tests refer to those with the same input but with expected outputs that vary depending on which intent the program satisfies. Given the initial set of tests, adversarial prompting directs the LLM to “Modify the expected output according to the adversarial function intent” (P8). This allows us to align all test oracles to the same inputs and count the number of differing outputs. Inspired by mutation testing (jia2010analysis-mutation, ), we measure the degree of adversarial conflict in each intent with the score $adversarial\_score=\frac{different\_tests}{all\_tests}$. We define an adversarial threshold as $thres=\frac{100\%}{K}$, where K is the number of inferred intents. For example, in Figure 2, $intents\_1$ and $intents\_2$ produce different outputs in two out of four generated tests, resulting in an $adversarial\_score$ of $50\%$.
@@ 3.3. Agentrepair
Agentrepair takes as input the program intents inferred by Agentreason and the tests generated by Agenttest. It aims to generate patches applicable to the input program that fix the bug by ensuring that both the original and adversarial test cases pass. The key difference between Agentrepair and prior conversation-based repair approaches is that Agentrepair explicitly explores the diversity and differences in the root causes of the bug, whereas prior work either limits themselves in terms of diversity by iterative refining based on a single previous patch (ChatRepair-ISSTA24, ; ITER, ) or increases randomness and cost by using random samples in patch generation through adjusting LLM temperatures (hidvégi2024cigar, ; kong2024contrastrepair, ).
@@
Next, based on these identified root causes, Agentrepair sends three separate prompts to guide the LLM in fixing the bug according to each root cause (P11, P12, and P13). Each program’s intent leads to three initial patches. However, not all correct patches can be generated in one attempt. Following prior work on conversational LLM approaches (ITER, ; ChatRepair-ISSTA24, ), Agentrepair refines previous patches by executing them against the test cases. If errors are found, the error messages are collected and provided to the LLM for further patch refinement (P14). The process is configured to allow a maximum of refinement configuration reaches.
3.4. Patch Validation
In patch validation, our approach introduces a novel evaluation of patch correctness using original tests, adversarial tests, and manual analysis.
First, we run all patches against the original test cases. Patches that pass these tests are considered plausible. Each plausible patch, generated from a specific inferred program intent, is expected to pass only the adversarial tests associated with that same intent. Assuming one of the inferred intents is correct, if a patch passes the adversarial tests for that intent, it is considered likely-correct with respect to that intent.
Finally, all likely-correct patches are manually analyzed by comparing them to the human-written patch. If judged to be semantically equivalent, they are labeled as believed-correct patches.
@@ 4. Evaluation
  • RQ4 (Cost): What is the cost of AdverIntent-Agent in generating plausible patches?
@@
Table 2. Evaluation Benchmark
@@
Defects4J 2.0 (Defects4J, ) 17 835
HumanEval-Java (jiang2023knod, ) 164 164
@@ 4.1.1. Benchmarks
We evaluate AdverIntent-Agent on the widely used benchmarks Defects4J 2.0 (Defects4J, ) and HumanEval-Java (jiang2023impact, ). Table 2 summarizes details of the considered benchmarks. Defects4J 2.0 consists of 835 bugs from 17 real-world projects and has been widely used in related work, enabling us to conduct a fair comparison. Additionally, we evaluate HumanEval-Java. In HumanEval-Java, developers converted Python programs and their associated test cases from HumanEval into Java programs and JUnit test cases, as well as some bugs were intentionally introduced into the correct Java programs. HumanEval-Java was released after the data collection used for training GPT-3.5, providing a new benchmark for evaluating the model’s ability to handle Java coding tasks.
4.1.2. Evaluating patch quality
@@
We measure APR effectiveness in terms of the number of both plausible (test-passing) patches and correct patches. Patches that overfit to the tests but fail to generalize to the desired specification are a well-known problem in automatic program repair (cure-worse-15, ; kali-issta15, ). We take multiple approaches to evaluate patch correctness. First, we assess whether the patch is plausible that makes the original test cases pass (kali-issta15, ). If a patch is plausible, we further identify whether it is an exact match to the developer-written ground truth patch with string comparison, ignoring whitespace differences.
For the remaining plausible patches that do not exactly match the ground truth, we employ an LLM to assess their correctness based on whether they pass one of the AI-generated tests. If a patch passes its intended test case, it is considered a likely-correct patch; otherwise, it is considered likely-overfitting. Finally, we manually assess patch correctness following prior work (ChatRepair-ISSTA24, ; acs, ; genpro2009, ; CURE-icse21, ). Two authors independently cross-check the correctness of these patches (CURE-icse21, ; ye2022selfapr, ; jiang2023knod, ; repairagent, ). If both authors agree that a patch is semantically equivalent to the ground truth, it is considered believed-correct. These, along with the exact match patches, form the set of correct patches.
@@ 4.1.3. Methodology
We compare AdverIntent-Agent with related work across different repair families, including fine-tuning-based methods: SelfAPR (ye2022selfapr, ), AlphaRepair (alpha-repair-fse22, ), ITER (ITER, ); conversation-based methods: ChatRepair (ChatRepair-ISSTA24, ), Cigar (hidvégi2024cigar, ), contrastRepair (kong2024contrastrepair, ), and a baseline from GPT-4; as well as the agent-based approach: RepairAgent (repairagent, ). We reimplemented experiments for ChatRepair (ChatRepair-ISSTA24, ) based on GPT-4 and also conducted a baseline experiment using GPT-4 without execution feedback. For the remaining approaches, we report quantitative results from the corresponding papers and repositories. Plausible and correct repair results are reported as discussed above. In the experiment assuming perfect fault localization, we follow prior work (repairagent, ; kong2024contrastrepair, ; ChatRepair-ISSTA24, ) by providing the buggy statements and their contextual code to the LLMs. For a more realistic comparison, we provide a complete buggy class tested by the failing test cases, allowing the LLM to achieve finer granularity at both the function and statement levels. Existing related work evaluated on HumanEval-Java is mostly based on LLM approaches. This is because HumanEval-Java was released recently, after the data collection used for training GPT-3.5, meaning that earlier approaches were not able to evaluate it at the time they published their paper.
@@
We study the effectiveness of adversarial reasoning by: 1) Measuring the adversarial degree between inferred intents to assess how different and distinct they are. This is based on the number of distinct test oracles generated from the same inputs. The evaluation is performed during the initial intent generation, without regenerating intents if the score falls below the threshold. 2) Comparing the alignment between inferred intents and ground-truth intents extracted from the human-written patch. We first use an LLM to generate the ground-truth intent from the patch, then assess semantic equivalence using the GPT-4o. 3) Evaluating its effectiveness as a fault localization and patch generation technique. For buggy statements spanning multiple locations, identifying any part of the buggy statement is considered correct, as it provides a strong starting point for patch generation. This is justified by prior work (ITER, ), which shows that subsequent buggy statements can be discovered iteratively. For patch generation, we compare the number of successful patches generated from the initial intent and the two adversarial intents to evaluate how adversarial reasoning promotes diversity. All three analyses involve substantial manual effort. To make this feasible, we conduct the study on 300 randomly selected bugs from Defects4J 2.0.
@@
In this research question, we are primarily concerned with the true positives—tests that successfully identify and filter out overfitting patches. RQ3 assesses the quality of test cases generated by Agenttest by evaluating their effectiveness in detecting likely-overfitting patches—that is, the extent to which the newly generated tests, as judged by GPT-4o, induce failures in plausible but incorrect patches. For all patches considered likely-overfitting by the LLM, we finally conduct manual analysis to confirm whether they are truly overfitting.
@@
In RQ4, we study the cost of AdverIntent-Agent in comparison to related work (ChatRepair-ISSTA24, ; hidvégi2024cigar, ; kong2024contrastrepair, ) that also make extensive use of LLM APIs. Specifically, we examine the token cost (i.e., the average number of tokens used to repair one bug), query cost (i.e., the average number of queries made to repair one bug), and ultimately, the estimated monetary cost.
Table 3. Comparison with State-of-the-art on three evaluation benchmarks: Defect4J 2.0 and HumanEval-Java. A * indicates our reimplementation results.
@@
Fine-tuning-based SelfAPR (ye2022selfapr, ) 130 110 107 67
AlphaRepair (alpha-repair-fse22, ) 109 79 90 50
ITER (ITER, ) - - 119 74
LLM Conversation-based ChatRepair (ChatRepair-ISSTA24, ) - 162 - -
Cigar (hidvégi2024cigar, ) - - 185 69
ContrastRepair (kong2024contrastrepair, ) 201 143 - -
Agent-based RepairAgent (repairagent, ) 186 164 - -
@@
ChatRepair * (ChatRepair-ISSTA24, ) 137 130 126 88
ContrastRepair (kong2024contrastrepair, ) 151 137 - -
Cigar (hidvégi2024cigar, ) - - 152 102
AdverIntent-Agent 155 140 146 105
@@ 4.2. RQ1: AdverIntent-Agent Effectiveness
Table 3 compares the effectiveness of AdverIntent-Agent with state-of-the-art in two considered benchmark Defects4J 2.0 and HumanEval-Java. The second column lists the existing APR approaches from three categories: tuning-based, LLM-based and Agent-based approaches. Their effectiveness results respectively based on perfect fault localization and realistic fault localization are given in the third and fourth column groups. The best performance number is highlighted in bold in the table. For example, in the last row of the Defects4J 2.0 benchmark, AdverIntent-Agent generated plausible patches for 135 bugs and correctly repaired 77 of them with realistic fault localization. These numbers increase to 180 and 141, respectively, when assuming perfect fault localization.
@@
Avoid overfitting with stronger patch specification by adversarial tests. AdverIntent-Agent achieves the highest number of correct repairs in both the Defects4J 2.0 benchmark (77 of 835 bugs repaired) and the HumanEval-Java benchmark (105 of 164 bugs repaired) under the realistic fault localization configuration. While AdverIntent-Agent does not produce the highest number of plausible patches, this is due to its stronger validation criteria: patches are assessed against both the original and adversarial test cases, where adversarial tests effectively filter out overfitting patches. In contrast, prior work only validates against original test cases, resulting in a higher number of overfitting patches that cannot be eliminated.
Figure 3 shows patches generated based on the same program intent for the add_elements program in HumanEval-Java, both with and without the use of dynamic, precise prompting to guide patch generation. Figure 3 provides an inferred program intent and identifies the faulty statements. There are two issues in this program. First, the buggy code considers at least two digits, which violates the program description of at most two digits (line 4). Second, the buggy code subtracts instead of adding to the final results. Both of these bugs need to be fixed. Figure 3 displays the initial patch generated by Agentrepair. Although both issues are addressed in the patch, it incorrectly identifies all numbers under 100 as having two digits, failing to account for negative numbers that may exceed two digits. The dynamic prompt, shown in Figure 3, first asks the LLM to identify the root cause of the patch to improve it. One response identifies the root cause as ”negative signs,” and a follow-up precise prompt guides the LLM to generate a patch that accounts for negative signs. This leads to a refined patch being generated. Based on the description, AdverIntent-Agent also generates an additional test case to expose the differing program behavior between the initial patch and the refined patch, where the expected program behavior is ’-4’, and the initial patch’s actual output is ’-16’.
@@
Rich information as output AdverIntent-Agent provides rich information, as shown in Figure 3, including inferred program intent and additional test cases, which offer valuable resources for developers to understand the patches. This is a significant distinction from prior work, which typically provides only patches. To our knowledge, AdverIntent-Agent is the only framework that integrates inferred program intents, tests, and patches. This paradigm shift enhances how developers engage with APR patches — enabling them from understanding the patch and choosing one program intent that most closely aligns with their original intent.
Figure 3. AdverIntent-Agent generates a correct patch based on dynamic explicit root cause guidance
@@
An inferred program intent and located faulty statements: Program Intent: Given an array of integers arr and an integer k, the program should return the sum of elements that have at most two digits, within the first k elements of arr. For this calculation, any negative sign should count as a digit, so negative numbers up to -99 (e.g., -5, -23) qualify as ”two-digit numbers” under this rule. Suspicious Faulty Statements - if( Integer.toString (arr.get(i)).length() ¿= 2) { //bug: at least instead of at most - result -= arr.get(i); //bug: Subtraction instead of addition - }
@@
//Bug: -10 will be incorrectly count
[…]
@@
[…]
Generate a patch to address negative signs.
[…]
Replace Math.abs(num)<100‘, convert the
number to a string and count characters.
((c)) Dynamic precise prompt to guide patch refinement
[…]
//Convert the number to a string
[…]
//Check at most two digits
[…]
//Addition instead of subtraction
[…]
((d)) A refined patch based on precise dynamic prompt in (c)
@Test
public void testNegativeNumbers() {
ArrayList arr = new ArrayList(Arrays.
asList(-1, -12, -3));
int k = 3;
// Expected output: -1 + (-3)
assertEquals(-4, add_elements(arr, k));
@@
Answer to RQ1: AdverIntent-Agent repairs the highest number of bugs, with 77 in Defects4J 2.0 and 105 in HumanEval-Java. This success is due to effective adversarial reasoning on program intents and dynamic, root-cause-specific prompt construction. The number of plausible yet overfitting patches is low, as all patches undergo robust validation against both original and adversarial tests. Additionally, AdverIntent-Agent offers comprehensive resources for function reasoning, bug analysis, and test cases, providing a unique package in program repair approaches. This design shifts the paradigm by enabling developers to understand APR patches and select program intents more effectively.
4.3. RQ2: Adversarial Intents Effectiveness
Table 4. Intent Alignment Statistics

### v2

@@ 3.2. Agenttest
First, Agenttest generates tests to measure the adversarial degree of inferred intents. If the adversarial degree of an intent is low, that intent will be removed, and a new adversarial intent will be generated. Second, Agenttest generates additional tests to address the overfitting problem by validating the patches generated by Agentrepair. Specifically, we validate whether the generated patches are overfitting to a specific intent. When one of the intents is close to the developer’s intent, the generated patches are considered correct rather than overfitting. Note that this use of tests generated as an integral part of repair is, to our knowledge, novel. Overfitting patch assessment tends to be considered a post-processing step in patch generation, or left to manual effort (as in previous agent-based approaches (Bouzenia et al., 2024; Lee et al., 2024; Zhang et al., 2024b), which conclude repair attempts upon finding plausible test-passing patches).
@@
Recall that Agentreason produces multiple adversarial program intents. Similarly, Agenttest aims to generate adversarial tests for these inferred intents. In this context, adversarial tests refer to those with the same input but with expected outputs that vary depending on which intent the program satisfies. Given the initial set of tests, adversarial prompting directs the LLM to “Modify the expected output according to the adversarial function intent” (P8). This allows us to align all test oracles to the same inputs and count the number of differing outputs. Inspired by mutation testing (Jia and Harman, 2010), we measure the degree of adversarial conflict in each intent with $adversarial\_score=\frac{different\_tests}{all\_tests}$. We define an adversarial threshold as $thres=\frac{100\%}{K}$, where K is the number of inferred intents. For example, in Figure 2, $intents\_1$ and $intents\_2$ produce different outputs in two out of four generated tests, resulting in an $adversarial\_score$ of $50\%$.
@@ 3.3. Agentrepair
Agentrepair takes as input the program intents inferred by Agentreason and the tests generated by Agenttest. It aims to generate patches applicable to the input program that fix the bug by ensuring that both the original and adversarial test cases pass. The key difference between Agentrepair and prior conversation-based repair approaches is that Agentrepair explicitly explores the diversity and differences in the root causes of the bug, whereas prior work either limits themselves in terms of diversity by iterative refining based on a single previous patch (Xia and Zhang, 2024; Ye and Monperrus, 2024) or increases randomness and cost by using random samples in patch generation through adjusting LLM temperatures (Hidvégi et al., 2024; Kong et al., 2024).
@@
Next, based on these identified root causes, Agentrepair sends three separate prompts to guide the LLM in fixing the bug according to each root cause (P11, P12, and P13). Each program’s intent leads to three initial patches. However, not all correct patches can be generated in one attempt. Following prior work on conversational LLM approaches (Ye and Monperrus, 2024; Xia and Zhang, 2024), Agentrepair refines previous patches by executing them against the test cases. If errors are found, the error messages are collected and provided to the LLM for further patch refinement (P14). The process is configured to allow a maximum of refinement configuration reaches.
@@ 4. Evaluation
  • RQ4 (Cost): What is the token cost of AdverIntent-Agent in generating plausible patches?
@@
Table 2. Evaluation benchmarks.
@@
Defects4J 2.0 (Just et al., 2014) 17 835
HumanEval-Java (Jiang et al., 2023b) 164 164
@@ 4.1.1. Benchmarks
We evaluate AdverIntent-Agent on the widely used benchmarks Defects4J 2.0 (Just et al., 2014) and HumanEval-Java (Jiang et al., 2023a). Table 2 summarizes details of the considered benchmarks. Defects4J 2.0 consists of 835 bugs from 17 real-world projects and has been widely used in related work, enabling us to conduct a fair comparison. Additionally, we evaluate HumanEval-Java. In HumanEval-Java, developers converted Python programs and their associated test cases from HumanEval into Java programs and JUnit test cases, as well as some bugs were intentionally introduced into the correct Java programs. HumanEval-Java was released after the data collection used for training GPT-3.5, providing a new benchmark for evaluating the model’s ability to handle Java coding tasks.
4.1.2. Patch Quality Evaluation
@@
We measure APR effectiveness in terms of the number of both plausible (test-passing) patches and correct patches. Patches that overfit to the tests but fail to generalize to the desired specification are a well-known problem in automatic program repair (Smith et al., 2015; Qi et al., 2015). We take multiple approaches to evaluate patch correctness. First, we assess whether the patch is plausible that makes the original test cases pass (Qi et al., 2015). If a patch is plausible, we further identify whether it is an exact match to the developer-written ground truth patch with string comparison, ignoring whitespace differences.
For the remaining plausible patches that do not exactly match the ground truth, we employ an LLM to assess their correctness based on whether they pass LLM-generated tests for based on one specific program intent. If a patch passes its intended test case, it is considered a likely-correct patch; otherwise, it is considered likely-overfitting. Finally, we manually assess patch correctness following prior work (Xia and Zhang, 2024; Xiong et al., 2017; Weimer et al., 2009; Jiang et al., 2021). Two authors independently cross-check the correctness of these patches (Jiang et al., 2021; Ye et al., 2022b; Jiang et al., 2023b; Bouzenia et al., 2024). If both authors agree that a patch is semantically equivalent to the ground truth, it is considered believed-correct. These, along with the exact match patches, form the set of correct patches.
@@ 4.1.3. Methodology
We discuss our methodology for each research question below.
@@
We compare AdverIntent-Agent with related work across different repair families, including fine-tuning-based methods: SelfAPR (Ye et al., 2022b), AlphaRepair (Xia and Zhang, 2022), ITER (Ye and Monperrus, 2024); conversation-based methods: ChatRepair (Xia and Zhang, 2024), Cigar (Hidvégi et al., 2024), ContrastRepair (Kong et al., 2024), and a baseline from GPT-4; as well as the agent-based approach: RepairAgent (Bouzenia et al., 2024). We reimplemented experiments for ChatRepair (Xia and Zhang, 2024) based on GPT-4 and also conducted a baseline experiment using GPT-4 without execution feedback. For the remaining approaches, we report quantitative results from the corresponding papers and repositories. Plausible and correct repair results are reported as discussed above. In the experiment assuming perfect fault localization, we follow prior work (Bouzenia et al., 2024; Kong et al., 2024; Xia and Zhang, 2024) by providing the buggy statements and their contextual code to the LLMs. For a more realistic comparison, we provide a complete buggy class tested by the failing test cases, allowing the LLM to achieve finer granularity at both the function and statement levels. Existing related work evaluated on HumanEval-Java is mostly based on LLM approaches. This is because HumanEval-Java was released recently, after the data collection used for training GPT-3.5, meaning that earlier approaches were not able to evaluate it at the time they published their paper.
@@
We study the effectiveness of adversarial reasoning by: 1) Measuring the adversarial degree between inferred intents to assess how different and distinct they are. This is based on the number of distinct test oracles generated from the same inputs. The evaluation is performed during the initial intent generation, without regenerating intents if the score falls below the threshold. 2) Comparing the alignment between inferred intents and ground-truth intents extracted from the human-written patch. We first use an LLM to generate the ground-truth intent from the patch, then assess semantic equivalence using the GPT-4o. 3) Evaluating its effectiveness as a fault localization and patch generation technique. For buggy statements spanning multiple locations, identifying any part of the buggy statement is considered correct, as it provides a strong starting point for patch generation. This is justified by prior work (Ye and Monperrus, 2024), which shows that subsequent buggy statements can be discovered iteratively. For patch generation, we compare the number of successful patches generated from the initial intent and the two adversarial intents to evaluate how adversarial reasoning promotes diversity. All three analyses involve substantial manual effort. To make this feasible, we conduct the study on 300 randomly selected bugs from Defects4J 2.0.
@@
In this research question, we explore to what extent newly generated tests can successfully identify and filter out overfitting patches. To evaluate this, we adopt a two-step approach that combines LLM-based selection with manual confirmation. Specifically, we employ GPT-4o, which takes as input a set of plausible patches and a set of newly generated tests. First, we perform a sanity check on the generated tests by assessing their correctness and removing those with low-confidence oracles. Next, we rank the tests and select the top 70% from the ranked list. These selected tests are then executed against the plausible patches to identify likely overfitting patches, defined as those that fail the tests. For all patches flagged as likely overfitting by the LLM, we conduct a final manual analysis to determine the actual number of true positives.
@@
In RQ4, we conduct a detailed analysis of the cost associated with AdverIntent-Agent by comparing it to several related methods, including ChatRepair(Xia and Zhang, 2024), Cigar (Hidvégi et al., 2024), and RepairAgent (Bouzenia et al., 2024), all of which rely heavily on large language model (LLM) APIs for automated program repair tasks. Our comparison focuses specifically on the token cost, which refers to the average number of tokens consumed by each approach to repair a bug in the Defect4J benchmark. We evaluate the token cost in real-world software projects, thus we do not evaluate the token cost on HumanEval-Java. By evaluating token usage in this context, we aim to provide a clear understanding of the computational resources required by each method and to highlight the efficiency and practicality of AdverIntent-Agent in real-world software engineering scenarios.
Table 3. Comparison with state-of-the-art on three evaluation benchmarks: Defect4J 2.0 and HumanEval-Java. A ‘*’ indicates our reimplementation results, and a ‘–’ indicates that the result is not available in the literature. The best performance is shown in bold.
@@
Fine-tuning-based SelfAPR (Ye et al., 2022b) 130 110 107 67
AlphaRepair (Xia and Zhang, 2022) 109 79 90 50
ITER (Ye and Monperrus, 2024) - - 119 74
LLM Conversation-based ChatRepair (Xia and Zhang, 2024) - 162 - -
Cigar (Hidvégi et al., 2024) - - 185 69
ContrastRepair (Kong et al., 2024) 201 143 - -
Agent-based RepairAgent (Bouzenia et al., 2024) 186 164 - -
@@
ChatRepair * (Xia and Zhang, 2024) 137 130 126 88
ContrastRepair (Kong et al., 2024) 151 137 - -
Cigar (Hidvégi et al., 2024) - - 152 102
AdverIntent-Agent 154 140 146 105
@@ 4.2. RQ1: AdverIntent-Agent Effectiveness
Table 3 compares the effectiveness of AdverIntent-Agent with state-of-the-art in two considered benchmark Defects4J 2.0 and HumanEval-Java. The second column lists the existing APR approaches from three categories: fine-tuning-based, LLM-based and Agent-based approaches. Their effectiveness results respectively based on perfect fault localization and realistic fault localization are given in the third and fourth column groups. The best performance number is highlighted in bold in the table. For example, in the last row of the Defects4J 2.0 benchmark, AdverIntent-Agent generated plausible patches for 135 bugs and correctly repaired 77 of them with realistic fault localization.
AdverIntent-Agent achieves the highest number of correct repairs in both the Defects4J 2.0 benchmark (77 of 835 bugs repaired) and the HumanEval-Java benchmark (105 of 164 bugs repaired) under the realistic fault localization configuration. For Defects4J 2.0 benchmark, these numbers increase to 180 and 141, respectively, when assuming perfect fault localization is known. For HumanEval-Java benchmark, these numbers increase to 154 and 140, respectively, when assuming perfect fault localization is known. Our result shows the overall effectiveness of AdverIntent-Agent.
@@
While AdverIntent-Agent does not produce the highest number of plausible patches, this is due to its stronger validation criteria: patches are assessed against both the original and adversarial test cases, where adversarial tests effectively filter out overfitting patches. In contrast, prior work only validates against original test cases, resulting in a higher number of overfitting patches that cannot be eliminated.
We present a case study to demonstrate the effectiveness of AdverIntent-Agent. Figure 3 shows patches generated based on the same program intent for the add_elements program in HumanEval-Java. Figure 3 provides an initial inferred program intent and identifies the faulty statement in which the given code checks for at least two digits, which violates the inferred intent of at most two digits (see the comment in Figure 3).
Figure 3 shows the initial patch generated by Agentrepair. Although this patch is intended to fix the issue, it still contains two bugs. First, the patch incorrectly identifies all numbers under 100 as having two digits, failing to account for negative numbers that may have more than two digits. Second, the initial intent and its patch fail to correct the buggy code that subtracts from, rather than adds to, the final result.
The adversarial intent, shown in Figure 3, critiques the initial intent and patch and infers a new intent that differs from the initial intent. This leads to the generation of an adversarial patch, shown in Figure 3. Based on the inferred program intent, AdverIntent-Agent also generates an additional test case Figure 3 to reveal the behavioral difference between the initial patch and the adversarial patch: the expected output is ’-4’ (making the adversarial patch pass), while the initial patch’s actual output is ’-16’ (which fails this test case).
@@
We can see AdverIntent-Agent provides rich information, including inferred program intent and newly generated test cases, which offer valuable resources for developers to understand the patches. This is a significant distinction from prior work, which typically provides only patches. To our knowledge, AdverIntent-Agent is the only framework that integrates inferred program intents, tests, and patches. This paradigm shift enhances how developers engage with APR patches, enabling them from understanding the patch and choosing one program intent that most closely aligns with their original intent.
@@
An inferred program intent and located faulty statements: Program Intent: Given an array of integers arr and an integer k, the program should return the sum of elements that have at most two digits, within the first k elements of arr. For this calculation, any negative sign should count as a digit, so negative numbers up to -99 (e.g., -5, -23) qualify as ”two-digit numbers” under this rule. Suspicious Faulty Statements - if( Integer.toString (arr.get(i)).length() ¿= 2) { //bug: at least instead of at most - result -= arr.get(i); - }
@@
An initial patch: //Bug: -10 will be incorrectly count - if( Integer.toString (arr.get(i)).length() ¿= 2) { + int num = arr.get(i); + if (Math.abs(num) ¡ 100) {
@@
An adversarial inferred program intent: //bug 1: at least instead of at most - if( Integer.toString (arr.get(i)).length() ¿= 2) { //bug 2: Subtraction instead of addition - result -= arr.get(i);
((c)) An adversarial inferred program intent and faulty statements
An adversarial patch: + int num = arr.get(i); + String numStr = Integer.toString(num); + int digitCount = numStr.length(); //Check at most two digits + if (digitCount ¡= 2) { + result += num; //Addition instead of subtraction + }
((d)) An adversarial patch
A generated test case: @Test public void testAddElements() { ArrayList arr = new ArrayList(Arrays. asList(-1, -12, -3)); int k = 3; // Expected output: -1 + (-3) assertEquals(-4, add_elements(arr, k)); }
@@
Figure 3. A correct patch generated by AdverIntent-Agent.
Answer to RQ1: AdverIntent-Agent repairs 77 in Defects4J and 105 in HumanEval-Java benchmarks. This success is due to effective adversarial reasoning on program intents and dynamic, root-cause-specific prompt construction. Additionally, AdverIntent-Agent offers comprehensive resources for function reasoning, bug analysis, and test cases, providing a unique package in program repair approaches. This design shifts the paradigm by enabling developers to understand APR patches and select program intents more effectively.
4.3. RQ2: Adversarial Intent Effectiveness
Table 4. Program intent alignment statistics.